OpenStudy (anonymous):

-1/15 x ? =1 solve please

Nnesha (nnesha):

\[\huge\rm \frac{ -1 }{ 15 }x=1\] first multiply both side by 15

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

Can u help with more?

Nnesha (nnesha):

what you got ??

OpenStudy (anonymous):

Which value is the product of −27 / 68 × −6, in simplified form?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

\[\frac{ -27 }{ 68 } \times -6\] ???

OpenStudy (anonymous):

Yes!

Nnesha (nnesha):

you just have to multiply numerator by numerator

OpenStudy (anonymous):

So... \[-27 x -6\]

Nnesha (nnesha):

nope is that x ??

OpenStudy (anonymous):

times yeah

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

so multiply

OpenStudy (anonymous):

-27 x -6 ?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Okay! more?

Nnesha (nnesha):

then reduce the fraction

OpenStudy (anonymous):

Okay... dont know how

OpenStudy (anonymous):

@Godlovesme @Nnesha

Nnesha (nnesha):

-27 times -6 = ??

OpenStudy (anonymous):

162?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Godlovesme

Nnesha (nnesha):

 now reduce this fraction